OpenStudy (anonymous):

7 over 35 = 28 over ?

OpenStudy (anonymous):

can't wait to see this

Parth (parthkohli):

I would like to work with the relation. the denominator is 5 times the numerator.

Parth (parthkohli):

simply multiply 28 with 5.

Parth (parthkohli):

Do you know multiplication?

OpenStudy (anonymous):

\[\frac{7}{35}\times \frac{4}{4}=\frac{28}{?}\]

OpenStudy (anonymous):

no nothing wrong with any method

OpenStudy (diyadiya):

:D